# Strongly Interacting Dark Matter admixed Neutron Stars

### v2

@@ Abstract
Dark matter may accumulate in neutron stars given its gravitational interaction and abundance. We investigate the modification of neutron star properties and confront them with the observations in the context of strongly-interacting dark matter scenario, specifically for a QCD-like theory with G2 gauge group for which a first-principles equation-of-state from lattice calculations is available. We study the impact of various observational constraints and modeling of the QCD equation of state on the combined neutron stars. The results indicate that dark matter masses of a few hundred MeV to a few GeV are consistent with the latest observed neutron star properties.
@@ 1 Introduction
The dark matter hypothesis Rubin:1978kmz ; Bambi:2015mba ; Cirelli:2024ssz is and remains one of the primary candidates to resolve a multitude of astronomical observations, despite a lack of direct observation. Assuming the existence of dark matter, its gravitational interactions are a given, with only the masses and species composition, if any, as inputs. In addition, indirect evidence supports a sizable dark matter self interaction Spergel:1999mh ; Kaplinghat:2015aga ; Tulin:2017ara ; Sagunski:2020spe ; Andrade:2020lqq ; Eckert:2022qia ; Bose:2022obr ; Adhikari:2022sbh ; Salucci:2018hqu , while interactions with standard model matter (in the following called ordinary) remain only bounded from above Bambi:2015mba ; Cirelli:2024ssz . Concentrating on dark matter gravitational signals is therefore a prudent strategy. However, except for supermassive dark matter particles Carr:2020gox , individual dark matter particles will not create an observable gravitational signal. Thus, only accumulations of dark matter particles will provide relevant signals, requiring an accumulator.
@@
Particle dark matter will necessarily accumulate in neutron stars making them an excellent laboratory to study dark matter gravitationally on small scales Akmal:1998cf ; Lattimer:2000nx ; Ozel:2016oaf . The amount of dark matter may vary depending on the age, history, and location of the neutron star, but most estimates conclude that that it is very unlikely to exceed 1% of the total neutron star mass Guver:2012ba ; Biesdorf:2024dor ; Hajkarim:2024ecp ; Kunkel:2024otq ; Wystub:2021qrn ; Pitz:2024xvh ; Diedrichs:2023trk ; Dengler:2021qcq ; Tolos:2015qra ; Mukhopadhyay:2015xhs ; Rezaei:2016zje ; Panotopoulos:2017idn ; Nelson:2018xtr ; Ellis:2018bkr ; Gresham:2018rqo ; Ivanytskyi:2019wxd ; Karkevandi:2021ygv ; Sen:2021wev ; Guha:2021njn ; Goldman:2013qla ; Xiang:2013xwa ; Li:2012ii ; Sandin:2008db ; Leung:2011zz ; Leung:2012vea ; Barbat:2024yvi ; Hippert:2021fch ; Issifu:2024htq ; Karkevandi:2024vov ; Shakeri:2022dwg ; Ema:2024wqr . Nevertheless, a careful analysis of potential effects of dark matter on gravitational and electromagnetic signals from neutron stars is necessary to assess if, and if yes what, impact dark matter could have on its properties. This will allow to understand whether neutron stars can be used to learn about dark matter. It also will clarify whether dark matter could, potentially at least, create effects relevant to understanding the ordinary matter aspects of neutron stars. Both aspects require a good data basis on neutron stars.
@@
On the observational side, fortunately, we do have a very good overview of global neutron star properties from observations. Among the direct observations, the Neutron star Interior Composition ExploreR (NICER) mission is dedicated to direct observation of neutron star properties via observations in the soft (0.2–12 keV) X-ray band Gendreau2024 . The mission concentrates on measurements of neutron star mass and radius, and will provide in its runtime a wide range of characterized neutron stars. Thus, a stochastically large enough sample is available to also detect variations of neutron star properties due to varying amounts of dark matter admixtures. In contrast to a single-component neutron star of ordinary matter, which yields a unique mass-radius relation, differing amounts of dark matter will fuzzy out such a curve. Thus, first-principle knowledge of this fuzzying, which we will present here, will allow to detect dark matter accumulation.
@@
Moreover, neutron stars with a large amount of dark matter in the interior, also known as mixed neutron stars, might provide smoking gun signals like supplementary peaks in the gravitational wave power spectral density Giangrandi:2022wht ; Ellis:2017jgp . Mixed neutron star binaries are thus a further testbed for dark matter properties, as they are capable of generating exotic signals for gravitational waves experiments. In particular, the so-called tidal deformability, indicating the deformation of colliding neutron stars, has been constrained by LIGO observations of a single binary neutron star merger GW170817 LIGOScientific:2017vwq ; LIGOScientific:2017ync . Along with this LIGO has also measured the “chirp mass” for the same system. These observations have been combined with simultaneous observations from other multi-messenger telescopes, leading to an even more comprehensive analysis LIGOScientific:2017ync .
@@
On the theoretical side, a primary ingredient in understanding such observable effects is the equation of state, which determines the evolution of thermodynamic properties, for both the ordinary component and a potential dark matter component inside the neutron star. Owing to the sign problem Gattringer:2010zz , lattice QCD, – the methodological mainstay to determine much of QCD physics – is unable to provide a fully controlled equation of state of ordinary matter inside neutron stars. Limited knowledge can be derived from heavy-ion experiments, models, and other methods Friman:2011zz . However, this has so far been insufficient to provide a complete understanding. In fact, observed neutron star properties are used to constrain the QCD equation of state Akmal:1998cf ; Lattimer:2000nx ; Ozel:2016oaf ; LIGOScientific:2017ync , under the assumption that neutron stars are primarily composed of ordinary matter. As a consequence, we will employ here three different candidates for the equation of state of ordinary matter, to represent this uncertainty. The particular choices will be discussed in section 2. However, they are made to reduce model dependence as much as possible in favor of data-driven input.
@@
Therefore, while neutron stars may be efficient dark matter accumulators, and consequently dark matter discovery avenues, insufficient knowledge on the accumulated amount coupled with no full control over neutron star ordinary physics present a challenging situation for the subject of dark matter admixed neutron stars. Progress with this conundrum would therefore be made if at least the dark matter equation of state would be controlled from first principles, reducing the uncertainties on the evolution of one of the two neutron star components. This is possible for WIMPs (and similar) cases, and several studies have been performed in this regard Quddus:2019ghy ; Quddus:2020lbt .
@@
Owing to the possible existence of large dark matter self-interactions Spergel:1999mh ; Kaplinghat:2015aga ; Tulin:2017ara ; Sagunski:2020spe ; Andrade:2020lqq ; Eckert:2022qia ; Bose:2022obr ; Adhikari:2022sbh ; Salucci:2018hqu , QCD-like hidden sectors have recently been in the spotlight as the resulting dark hadrons can inherently generate the required self-interactions. Such dark sectors thus demand their own equation of state calculations, should neutron stars be used to probe their existence. Similar to ordinary matter, these calculations may, however, also suffer from a sign problem. Fortunately, for some of the cases, like the Sp($N$)-based QCD-like theories  Hochberg:2014dra ; Hochberg:2014kqa ; Kulkarni:2022bvh , the sign problem is absent, and calculations may be possible, or have been done like for a dark matter model based on a two-color QCD-like hidden sector Drach:2015epq , see, e. g., Cotter:2012tt ; Boz:2019enj ; Hands:2024nkx .
@@
But, both, Sp($N$)-based and two-color based QCD-like strongly-interacting dark matter scenarios, feature only bosonic hadrons. Such dark matter systems do not experience stabilization by Fermi pressure, and thus gravitational collapse will only be halted when the dark quark substructure becomes relevant. This happens usually only at asymptotically high densities Friman:2011zz , yielding rather a dense dark matter nugget at the center of the neutron star at best, unless strong self-interactions provide stabilization Liebling:2012fv .
@@
Therefore any study of mixed neutron stars involving a new QCD-like admixture will benefit from a dark QCD-like sector with fermionic dark hadrons, where the equation of state is accessible in first-principle calculations. Fortunately, such possibilities exist, e. g. , a hidden QCD-like sector based on the gauge group G2 Holland:2003jy ; Maas:2012wr ; Wellegehausen:2013cya . It has already been shown that this theory can support pure compact stellar objects Hajizadeh:2017jsw , which could be interpreted as pure dark stars. Already the one-flavor theory provides a conserved quantum number carried by the fermionic dark baryons Maas:2012wr , and thereby offers a dark matter candidate, alongside a rich spectrum of further states Wellegehausen:2013cya . The two-flavor case furthermore offers dark pions as additional stable states with a Wess-Zumino-Witten-like $3\to 2$ process, relevant to the dark matter relic density generation mechanism Hochberg:2014dra ; Hochberg:2014kqa ; Kulkarni:2022bvh . There is also an additional $3\to 2$ process which converts dark pions into dark fermionic baryons, and allows to build Fermi pressure, from accrued bosonic dark matter, required for neutron star stabilization Hajizadeh:2017jsw . G2-QCD therefore provides an ideal testbed for the previously mentioned challenge.
@@
In this work, as a proof-of-principle, we use the available Maas:2012wr ; Wellegehausen:2013cya ; Hajizadeh:2017jsw equation of state for a dark G2-QCD sector with one flavor [^1] Since the chiral dynamics is the same for more flavors Holland:2003jy ; Wellegehausen:2013cya ; Kogut:2000ek , and up to multiplicities also the spectrum, we do not expect too may non-trivial changes to the equation of state when increasing the number of flavors slightly., and build neutron stars made from both dark matter and ordinary matter. This equation of state has been obtained using lattice simulations Maas:2012wr ; Wellegehausen:2013cya , and thus fully non-perturbatively. However, it lacks the quantitative precision of ordinary equations of state obtained indirectly, due to limitations in terms of computing time Wellegehausen:2013cya . Nonetheless, as already seen in the studies of pure dark objects Hajizadeh:2017jsw the accuracy should be more than sufficient for an exploratory study of the viability of such a scenario.
@@
For ordinary matter, we use a selection of the currently favored equations of state Kurkela:2014vha ; Dengler:2021qcq ; Pitz:2024xvh covering a wide range of phenomenology, and as much as possible data-driven constraints. This allows us to minimize the model-dependency of our investigation. The results will therefore only be quantitatively relevant if either we keep the amount of dark matter limited such that, within errors, neutron star properties remain the same. Or we stipulate that most neutron stars do not contain large amounts of dark matter, but some do, giving a distinct signal differing from ordinary neutron stars when observed. Alternatively, one can interpret our injection of G2-QCD equation of state seriously and demand modification of the ordinary matter equation of state to compensate for large modifications in the neutron star properties. However, this is beyond the scope of the present exploratory work, especially in view of the restricted set of available equations of state for the G2-QCD case. We will also assume throughout that the interactions between dark matter and ordinary matter, even at large densities of both, are negligible. This may be justified given that in most strongly-interacting dark matter scenarios the portal couplings to the SM are highly suppressed Hochberg:2015vrg ; Berlin:2018tvf ; Pomper:2024otb ; Bernreuther:2023kcg .
@@
For the sake of completeness, we discuss briefly our choices of the equations of state in section 2, together with a brief overview of the salient features of G2-QCD. Our approach to building a two-component neutron star is standard Hajkarim:2024ecp , and briefly rehearsed in section 3. In section 4 we provide results for a range of dark matter masses and dark matter content. We provide the mass-radius relations, together with maximal and minimal values for both. We also provide the tidal deformability, which is the primary input to the gravitational wave signal for the merging of two neutron stars. Finally, taking our inputs at face value, we interpret our result in terms of limits on dark matter properties or dark matter contents in neutron stars. This part should be considered rather a blueprint than a hard statement, given our limited inputs, but can be improved arbitrarily by improving these inputs. Additional results are available in appendices 6.2 and 6.3.
@@ 2 Equations of state
Figure 1: The equations of state from Kurkela:2014vha and Wellegehausen:2013cya used in this work. The dark matter candidate mass is chosen to be the mass of the neutron for this plot for comparison. The end of the lines indicate the maximum value for the pressure for the respective equation of state used in this work. The black line indicates $\varepsilon=p$. For the ordinary matter equation of state, the stars denote the results from NChPT while the dots show the endpoints of the piecewise polytropes.
The equation of state is the key ingredient for the Tolman-Oppenheimer-Volkoff (TOV) equation Oppenheimer:1939ne which are commonly used to calculate macroscopic neutron star observables. To describe a neutron star consisting of two fluids we need the equations of state of both of them. A novelty of this work is the use of an equation of state of a confining, asymptotically free theory obtained from first principles. This is crucial as strongly-interacting dark matter models make use of gauge theories often without investigating its non-perturbative signatures. We want to complement this by the use of equations of states for ordinary matter in as much a model agnostic way that is able to meet the observational constraints. In this way we minimize the number of free parameters. We emphasize that this choice is not sufficient to cover all possible scenarios for neutron stars, as equations of states with (first order) phase transitions or twin-stars, for example, are not covered Mendes:2024hbn ; Christian:2023hez .
For both ordinary and dark matter, we will use polytropes to interpolate between data points. In Kurkela:2009gj , this was done to interpolate between different physical regimes. We use the same method to interpolate between data points, also for both the interpolation of the NChPT data points and the lattice data points. Details of this procedure are given in appendix 6.1.

### v3

@@ Abstract
Dark matter may accumulate in neutron stars given its gravitational interaction and abundance. We investigate the influence of strongly-interacting dark matter, described by a QCD-like one-flavor $G_{2}$ gauge theory, on neutron stars. This choice allows to test, for the first time, a first-principles-determined non-Abelian dark matter equation of state, which supports composite fermionic dark matter and thus a Fermi-pressure-stabilized dark matter component. The ordinary matter part of the mixed star is described by available model-agnostic equations of state that interpolate between the low-density regime and high-density regime. We find that strongly-interacting dark matter has a similar impact on neutron stars as other model equation of states and confirm that strongly-interacting dark matter can be accommodated by constraints from neutron star observations within our uncertainties.
@@ 1 Introduction
The dark matter hypothesis Rubin et al. (1978); Bambi and Dolgov (2015); Cirelli et al. (2024) is and remains one of the primary candidates to resolve a multitude of astronomical observations, despite a lack of direct observation. Assuming the existence of dark matter, it interacts gravitationally by definition. In addition, indirect evidence is compatible with a sizable dark matter self-interaction Spergel and Steinhardt (2000); Kaplinghat et al. (2016); Tulin and Yu (2018); Sagunski et al. (2021); Andrade et al. (2021); Eckert et al. (2022); Bose and others (2022); Adhikari and others (2022); Salucci (2019), while interactions with standard model matter (in the following called ordinary) remain only bounded from above Bambi and Dolgov (2015); Cirelli et al. (2024). Concentrating on dark matter gravitational signals is therefore a prudent strategy. However, except for supermassive dark matter particles Carr et al. (2021), individual dark matter particles will not create an observable gravitational signal. Thus, only accumulations of dark matter particles will provide relevant signals, requiring an accumulator.
@@
Particle dark matter will necessarily accumulate in neutron stars making them an excellent laboratory to study dark matter gravitationally on small scales Akmal et al. (1998); Lattimer and Prakash (2001); Özel and Freire (2016). The possible amount of dark matter may vary depending on the age, history, and location of the neutron star, but most estimates conclude that that it is very unlikely to exceed 1% of the total neutron star mass Güver et al. (2014); Giangrandi et al. (2023, 2025), though numbers as large as up to 10% have also been motivated Grippa et al. (2025); Del Popolo et al. (2020). Nevertheless, a careful analysis of potential effects of dark matter on gravitational and electromagnetic signals from neutron stars is necessary to assess if, and if yes what, impact dark matter could have on its properties. This will allow to understand whether neutron stars can be used to learn about dark matter. It also will clarify whether dark matter could, potentially at least, create effects relevant to understanding the ordinary matter aspects of neutron stars. Both aspects require a good data basis on neutron stars.
@@
On the observational side, fortunately, we do have a very good overview of global neutron star properties from observations. Among the direct observations, the Neutron star Interior Composition ExploreR (NICER) mission is dedicated to direct observation of neutron star properties via observations in the soft (0.2–12 keV) X-ray band Gendreau et al. (2024). The mission concentrates on measurements of neutron star mass and radius, and will in its runtime provide a wide range of characterized neutron stars. Thus, a stochastically large enough sample is available to also detect variations of neutron star properties due to varying amounts of dark matter admixtures. In contrast to a single-component neutron star of ordinary matter, which yields a unique [^1] There can, of course, be variation due to temperature (and thus age), chemical composition in the crust and atmosphere, and further effects, also fuzzing out the mass-radius relation of pure ordinary matter neutron stars. These effects can, at least in principle, be quantified and accounted for. Any additional dark matter components will add on these effects, and will yield a deviation. mass-radius relation, differing amounts of dark matter will fuzzy out such a curve. Thus, first-principle knowledge of this fuzzying will allow to detect dark matter accumulation.
@@
Moreover, neutron stars with a large amount of dark matter in the interior, also known as mixed neutron stars, might provide smoking gun signals like supplementary peaks in the gravitational wave power spectral density Giangrandi et al. (2023); Ellis et al. (2018a). Mixed neutron star binaries are thus a further testbed for dark matter properties, as they are capable of generating exotic signals for gravitational waves experiments. In particular, the so-called tidal deformability, indicating the deformation of colliding neutron stars, has been constrained by LIGO observations of a single binary neutron star merger GW170817 Abbott and others (2017a, b). Along with this LIGO has also measured the “chirp mass” for the same system. These observations have been combined with simultaneous observations from other multi-messenger telescopes, leading to an even more comprehensive analysis Abbott and others (2017b).
@@
On the theoretical side, a primary ingredient in understanding such observable effects is the equation of state, which determines the evolution of thermodynamic properties, for both the ordinary component and a potential dark matter component inside the neutron star. Owing to the sign problem Gattringer and Lang (2010), lattice QCD – the methodological mainstay to determine much of QCD physics – is unable to provide a fully controlled equation of state of ordinary matter inside neutron stars. Limited knowledge can be derived from heavy-ion experiments, models, and other methods Friman et al. (2011). However, this has so far been insufficient to provide a final and complete understanding. In fact, observed neutron star properties are used to constrain the QCD equation of state Akmal et al. (1998); Lattimer and Prakash (2001); Özel and Freire (2016); Abbott and others (2017b), under the assumption that neutron stars are primarily composed of ordinary matter. Therefore, we will use three different candidates for the equation of state of ordinary matter to represent this uncertainty. The particular choices will be discussed in Section˜2. However, they are made to reduce model dependence as much as possible in favor of data-driven input.
@@
Thus, while neutron stars may accumulate some amount of dark matter, and consequently could be dark matter discovery avenues, insufficient knowledge on the accumulated amount coupled with remaining uncertainties over neutron star ordinary physics present a challenging situation for the subject of dark matter admixed neutron stars. Progress with this conundrum would therefore be made if at least the dark matter equation of state would be controlled from first principles, reducing the uncertainties on the evolution of one of the two neutron star components. This is possible for weakly interacting massive particles (and similar) cases, and several studies have been performed in this regard Quddus et al. (2020b, a). There have also been numerous other studies, which have been based on various assumptions and simplifications of the details of the dark matter equation of state Güver et al. (2014); Biesdorf et al. (2024); Hajkarim et al. (2024); Kunkel et al. (2024); Wystub et al. (2023); Pitz and Schaffner-Bielich (2024); Diedrichs et al. (2023); Dengler et al. (2022); Tolos et al. (2015); Mukhopadhyay and Schaffner-Bielich (2016); Rezaei (2017); Panotopoulos and Lopes (2017); Nelson et al. (2019); Ellis et al. (2018b); Gresham and Zurek (2019); Ivanytskyi et al. (2020); Karkevandi et al. (2022); Sen and Guha (2021); Guha and Sen (2021); Goldman et al. (2013); Xiang et al. (2014); Li et al. (2012); Sandin and Ciarcelluti (2009); Leung et al. (2011, 2012); Barbat et al. (2024); Hippert et al. (2022); Issifu et al. (2024); Karkevandi et al. (2024); Shakeri and Karkevandi (2024); Ema et al. (2025).
@@
Owing to the possible existence of large dark matter self-interactions Spergel and Steinhardt (2000); Kaplinghat et al. (2016); Tulin and Yu (2018); Sagunski et al. (2021); Andrade et al. (2021); Eckert et al. (2022); Bose and others (2022); Adhikari and others (2022); Salucci (2019), QCD-like hidden sectors have recently been in the spotlight as the resulting dark hadrons can inherently generate the required self-interactions. Such dark sectors thus demand their own equation of state calculations, should neutron stars be used to probe their existence. Similar to ordinary matter, these calculations may, however, also suffer from a sign problem. Fortunately, for some of the cases, like the Sp($N$)-based QCD-like theories  Hochberg et al. (2014, 2015); Kulkarni et al. (2023), the sign problem is absent, and calculations may be possible, or have been done like for a dark matter model based on a two-color QCD-like hidden sector Drach et al. (2016), see, e. g., Cotter et al. (2013); Boz et al. (2020); Hands et al. (2024).
@@
But, both, Sp($N$)-based and two-color based QCD-like strongly-interacting dark matter scenarios, feature only bosonic hadrons. Such dark matter systems do not experience stabilization by Fermi pressure, and thus gravitational collapse will only be halted when the dark quark substructure becomes relevant. This happens usually only at asymptotically high densities Friman et al. (2011), yielding rather a dense dark matter nugget at the center of the neutron star at best, unless strong self-interactions provide stabilization Liebling and Palenzuela (2023).
@@
Therefore any study of mixed neutron stars involving a new QCD-like admixture will benefit from a dark QCD-like sector with fermionic dark hadrons, where the equation of state is accessible in first-principle calculations. Fortunately, such possibilities exist, e. g. , a hidden QCD-like sector based on the gauge group G2 Holland et al. (2003); Maas et al. (2012); Wellegehausen et al. (2014). It has already been shown that this theory can support pure compact stellar objects Hajizadeh and Maas (2017), which could be interpreted as pure dark stars. Already the one-flavor theory provides a conserved quantum number carried by the fermionic dark baryons Maas et al. (2012), and thereby offers a dark matter candidate, alongside a rich spectrum of further states Wellegehausen et al. (2014). The two-flavor case furthermore offers dark pions as additional stable states with a Wess-Zumino-Witten-like $3\to 2$ process, relevant to the dark matter relic density generation mechanism Hochberg et al. (2014, 2015); Kulkarni et al. (2023). There is also an additional $3\to 2$ process, already in the one-flavor theory, which converts dark pions into dark fermionic baryons, and allows to build Fermi pressure, from accrued bosonic dark matter, required for neutron star stabilization Hajizadeh and Maas (2017). G2-QCD therefore provides an ideal testbed for the previously mentioned challenge.
@@
In this work, as a proof-of-principle, we use the available Maas et al. (2012); Wellegehausen et al. (2014); Hajizadeh and Maas (2017) equation of state for a dark G2-QCD sector with one flavor [^2] Since the chiral dynamics is the same for more flavors Holland et al. (2003); Wellegehausen et al. (2014); Kogut et al. (2000), and up to multiplicities also the spectrum, we do not expect too many non-trivial changes to the equation of state when increasing the number of flavors., and build neutron stars made from both dark matter and ordinary matter. This equation of state has been obtained using lattice simulations Maas et al. (2012); Wellegehausen et al. (2014), and thus fully non-perturbatively. However, it lacks the quantitative precision of ordinary equations of state obtained indirectly, due to limitations in terms of computing time Wellegehausen et al. (2014). Nonetheless, as already seen in the studies of pure dark objects Hajizadeh and Maas (2017) the accuracy should be more than sufficient for an exploratory study of the viability of such a scenario.
@@
For ordinary matter, we use a selection of the currently favored equations of state Kurkela et al. (2014); Dengler et al. (2022); Pitz and Schaffner-Bielich (2024) covering a wide range of phenomenology, and as much as possible data-driven constraints. This allows us to minimize the model-dependency of our investigation. The results will therefore only be quantitatively relevant if either we keep the amount of dark matter limited such that, within errors, neutron star properties remain the same. Or we stipulate that most neutron stars do not contain large amounts of dark matter, but some do, giving a distinct signal differing from ordinary neutron stars when observed. Alternatively, one can interpret our injection of G2-QCD equation of state seriously and demand modification of the ordinary matter equation of state to compensate for large modifications in the neutron star properties. However, this is beyond the scope of the present exploratory work, especially in view of the restricted set of available equations of state for the G2-QCD case. We will also assume throughout that the interactions between dark matter and ordinary matter, even at large densities of both, are negligible. This may be justified given that in most strongly-interacting dark matter scenarios the portal couplings to the SM are highly suppressed Hochberg et al. (2016); Berlin et al. (2018); Pomper and Kulkarni (2024); Bernreuther et al. (2024).
@@
For the sake of completeness, we discuss briefly our choices of the equations of state in Section˜2, together with a brief overview of the salient features of G2-QCD. Our approach to building a two-component neutron star is standard Hajkarim et al. (2024), and briefly rehearsed in Section˜3. In Section˜4 we provide results for a range of dark matter masses and dark matter content. We provide the mass-radius relations as well was the tidal deformability, which is the primary input to the gravitational wave signal for the merging of two neutron stars. Finally, taking our inputs at face value, we interpret our result in terms of limits on dark matter properties or dark matter contents in neutron stars. This part should be considered rather a blueprint than a hard statement, given our limited inputs, but can be improved arbitrarily by improving these inputs. Additional results are available in Section˜A.2.
@@ 2 Equations of state
Figure 1: The equations of state (left) and the speed of sound (right) from Kurkela et al. (2014) and Wellegehausen et al. (2014) used in this work. The dark matter candidate mass is chosen to be the mass of the neutron for this plot for comparison. The end of the lines indicate the maximum value for the pressure for the respective equation of state used in this work. The black line indicates $\varepsilon=p$ and $c_{s}^{2}=1$ respectively. For the ordinary matter equation of state, the stars denote the results from NChPT while the dots show the endpoints of the piecewise polytropes. For the dark matter equation of state, the points indicate lattice data with polytropes, see appendix Section˜A.1.
The equation of state (EoS) is the key ingredient for the Tolman-Oppenheimer-Volkoff (TOV) equation Oppenheimer and Volkoff (1939) which is commonly used to calculate macroscopic neutron star observables. To describe a neutron star consisting of two fluids we need the EoS of both of them. A novelty of this work is the use of an EoS of a confining, asymptotically free gauge theory obtained from first principles. The latter is crucial for a strongly-interacting dark matter model, as only this will ensure non-trivial constraints between different properties, e. g. the masses of stable states, resonances, and the EoS. We want to complement the dark matter EoS thus, naturally, by the use of an EoS for ordinary matter in as much a model agnostic way that is able to meet the observational constraints. In this way we minimize the number of free parameters. We emphasize that this choice is not sufficient to cover all possible scenarios for neutron stars, as equations of states with (first order) phase transitions or twin-stars, for example, are not covered Mendes et al. (2025); Christian et al. (2024).
Both, ordinary and dark matter, EoS will be needed to be interpolated from separate, partly discrete, information. This will be done using polytropes Kurkela et al. (2010). Details of this procedure are given in Section˜A.1.